Consider the general demand function:

Q1 = 8,000 - 16P + 0.75M + 30PR

a. Derive the equation for the demand function when M = $30,000 and P. = $50.

b. Interpret the intercept and slope parameters of the demand function derived in

part a.

C.

Sketch a graph of the demand function in part a. Where does the demand function

intersect the quantity-demanded axis? Where does it intersect the price axis?

d. Using the demand function from part a, calculate the quantity demanded when the

price of the good is $1,000 and when the price is $1,500.

e.

Derive the inverse of the demand function in part a. Using the inverse demand

function, calculate the demand price for 24,000 units of the good. Give an

interpretation of this demand price.


1
Expert's answer
2022-03-29T12:10:53-0400

Q1 = 8,000 - 16P + 0.75M + 30PR

i) Derive the equation for the demand function when M = $30,000 and P. = $50.

"Q1 = 8,000 - 16(50) + 0.75(30000)+ 30PR"

"Q1 = 8000-800+22500+30PR"

"Q1=29700+30PR"


b. Interpret the intercept and slope parameters of the demand function derived in

part a.

Q1=29700+30PR

This equation obeys y=mx +c

When x=0, The y intercept is "= 29700"

The slope from the equation is "= 30"


Using the demand function from part a, calculate the quantity demanded when the

price of the good is $1,000 and when the price is $1,500.

"Q1=29700+30(1000) = - 300"

"Q1=29700+30(1500)=- 15300"